# Data-driven Quantum Dynamical Embedding Method for Long-term Prediction on Near-term Quantum Computers

### v3

@@ Abstract
The increasing focus on long-term time series prediction across various fields has been significantly strengthened by advancements in quantum computation. In this paper, we introduce a data-driven method designed for long-term time series prediction with quantum dynamical embedding (QDE). This approach enables a trainable embedding of the data space into an extended state space, allowing for the recursive retrieval of time series information. Based on its independence of time series length, this method achieves depth-efficient quantum circuits that are crucial for near-term quantum computers. Numerical simulations demonstrate the model’s improved performance in prediction accuracy and resource efficiency over existing methods, as well as its effective denoising capabilities. We implement this model on the Origin “Wukong” superconducting quantum processor with a learnable error-cancellation layer (LECL) for error mitigation, further validating the practical applicability of our approach on near-term quantum devices. Furthermore, the theoretical analysis of the QDE’s dynamical properties and its universality enhances its potential for time series prediction. This study establishes a significant step towards the processing of long-term time series on near-term quantum computers, integrating data-driven learning with discrete dynamical embedding for enhanced forecasting capabilities.
@@ I Introduction
Figure 1: Circuits for QDE and Learning Protocols. (a) Illustration of a generic QDE architecture. The qubits for encoding the classical state are divided into two sets: the memory register (top, with $n_{m}$ qubits) for inputting memory $\bm{m}_{t}$, and the data register (bottom, with $n_{x}$ qubits) for inputting data $\bm{x}_{t}$ at time $t$. The circuit then undergoes evolution through a unitary operator $U(\bm{\theta})$, comprising parametrized quantum gates and fixed Hamiltonian evolution across $D$ layers. Finally, selected observable ensembles are measured to extract classical information. With this architecture and the initial point $(\bm{m}_{0},\bm{x}_{0})$, the state pairs $(\bm{m}_{1},\bm{x}_{1}),(\bm{m}_{2},\bm{x}_{2}),(\bm{m}_{3},\bm{x}_{3}),\dots$ are obtained in an autoregressive way. (b) The protocol of QRC and QRNNs. The data is injected into the quantum system and measured to retrieve classical information. The memory register maintains the quantum state form and will lose coherence in the long run. (c) An enhanced QDE architecture with $M$ linear-composite channels. In this design, the QDE is available to explore the dynamics with higher degrees of freedom. The final signal $\bm{x}_{t}$ is retrieved through the superposition of sub-signals $\bm{x}_{t}^{1},\dots,\bm{x}_{t}^{M}$ from different channels. (d) Training and predicting protocols. In the training stage, initial data $\bm{x}_{0}$ is fed into the QDE block, and the output data $\hat{\bm{x}}_{i}$ is used to compute the cost function, with $i={1,\cdots,L}$ and time updated over $L$ steps. After the training, the system undergoes an additional $T$ steps of evolution to predict future values. Note that while the memory evolves simultaneously with the data in both stages, it does not contribute to the cost function calculation.
@@
Recent years have witnessed a surge in the application of time series prediction across a range of fields, including finance, physics, and engineering, underscoring its significance [1, 2, 3]. Among the various types of predictions, long-term prediction is especially pivotal as it entails forecasting future values of a series that extend beyond the duration of the available historical data. This means that the prediction horizon is longer than the time span of the training dataset, presenting a unique challenge in identifying and extrapolating underlying patterns and trends that may continue to evolve over time [4, 5]. The importance of long-term prediction is evident in strategic planning for climate change, economic forecasting, and technological forecasting, where decisions made today are based on projections that reach far into the future [6, 7]. Classical algorithms for time series prediction predominantly rely on long short-term memory (LSTM) networks [8, 9], echo state networks (ESNs) [10, 11], and reservoir computing (RC) [12]. These data-driven methods have shown potential in capturing dynamical features. Recently, it has been demonstrated that classical data can be mapped into a quantum feature space using certain embeddings [13, 14, 15, 16], which is further utilized to approximate a variety of functions. These universal approximation properties [17, 18, 19, 20] facilitate the construction of inference functions for modeling temporal data patterns, suggesting the potential of quantum systems to effectively model the complex dynamics inherent in long-term time series.
@@
Current quantum algorithms for time series processing, such as quantum reservoir computing (QRC) [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], quantum recurrent neural networks (QRNNs) [33, 34, 35], and quantum dynamic mode decomposition (QDMD) [36], have demonstrated significant potential in understanding temporal dependencies in sequential data. However, QRC and QRNNs exhibit a linear increase in circuit depth as the length of the time sequence grows, posing challenges for noisy intermediate-scale quantum (NISQ) devices. The increasing circuit depth threatens the maintenance of quantum coherence over time, which is crucial for the accuracy of long-term predictions. Although the QRC model has shown significant performance of time series prediction, its requirement for deep circuits to maintain temporal dependencies makes it impractical for NISQ devices due to exacerbated decoherence and noise, which undermine the performance of quantum algorithms in realistic settings. Efficiently utilizing these devices for complex tasks like long-term time series prediction without deep circuits is essential for advancing practical quantum computing applications. Therefore, there is a pressing need for methods that can predict long-term time series with shallow quantum circuits, addressing the limitations posed by current quantum hardware. Solutions have been explored through both theoretical and physical implementations on various platforms [37, 38, 39, 40, 41, 42, 43], yet it remains an open question.
@@
In this study, we present a data-driven approach for time series prediction using quantum dynamical embedding (QDE). Specifically, we recurrently retrieve time series data from strong measurements of the QDE at the current time step. Our approach allows for the dynamic embedding of original data into an extended state space and facilitates the data-driven learning of embedding parameters. This method offers a solution for long-term time series prediction with circuit depth independent of the time series length and mitigates issues related to decoherence and error accumulation over time.
@@
We demonstrate the effectiveness of our method through numerical experiments with cosine-wave and composite signals. Additionally, we evaluate its denoising capabilities on the same prediction tasks when exposed to random noise, further showcasing the QDE’s versatility. Notably, a comparative analysis with existing methodologies reveals that our approach offers significant advantages in terms of average prediction accuracy and computational efficiency on the same tasks. This improvement highlights the robustness of our approach in capturing temporal dependencies more effectively than existing quantum models.
To validate the noise resistance of our method, we perform long-term predictions using a noisy simulator with a learnable error-cancellation layer (LECL) following the QDE. Additionally, we extend our implementation to the Origin “Wukong” superconducting quantum computer, integrating an LECL to generate time series. This demonstrates the potential of leveraging current quantum computing capabilities for real-world time series prediction challenges.
Theoretically, we explore the dynamics of a specific 2-qubit case, providing insights into the connection between the QDE and the inherent characteristics of time series. We present several theorems that underscore the universality of the QDE in this context. Finally, we adopt an autoregressive model in our approach because it allows us to utilize past observations to predict future values, which is a natural fit for time series data. The autoregressive framework aligns well with the principles of QDE, enabling efficient and accurate long-term predictions while maintaining the practical constraints of NISQ devices. Overall, this study advances our understanding of practical time series processing on near-term quantum devices, paving the way for more effective applications in quantum computing.
@@ II.1 The quantum dynamical embedding model
is the encoded input state. $U(\bm{\theta})$ is a trainable parameterized quantum circuit. The memory and data of the next step are extracted from observing of their corresponding subspaces, namely
@@
where $\mathcal{H}_{m}$ and $\mathcal{H}_{x}$ are the observable vector with length $n_{m}$ and $n_{x}$, respectively.
@@
The QDE distinguishes itself from previous studies of QRC and QRNNs in the manner of memory information storage. Unlike these models, where the quantum system serves as a memory reservoir for historical data, our approach overcomes the issue of coherence loss—a challenge in long-term data embedding and retrieval—through the QDE model, as illustrated in Fig. 1(b).
@@ II.2 Training and predicting
Building on the findings of previous research, the embedding of time series and the evolving of quantum systems require a circuit depth of $\mathcal{O}(L)$. Comparatively, our approach marks a notable advancement by achieving this with a significantly reduced circuit depth of $\mathcal{O}(\Gamma)$, where $\Gamma$ represents the base depth of a single QDE circuit block.
@@ II.3 Gradient evaluation
Referring to the definition of $\hat{x}_{t+1}$ in Eq. (3), this expression can be equivalently written as:
$\frac{\partial\hat{x}_{t+1}}{\partial\theta_{i}}$ $=\frac{\partial}{\partial\theta_{i}}\operatorname{Tr}\left(\mathcal{H}_{x}|\psi_{t}\rangle\langle\psi_{t}|\right).$ (6)
$=\frac{\partial}{\partial\theta_{i}}\mathrm{Tr}\left[\mathcal{H}_{x}U(\bm{\theta})U_{enc}(\hat{X}_{t})\rho_{0}U_{enc}^{\dagger}(\hat{X}_{t})U^{\dagger}(\bm{\theta})\right],$
@@
$\frac{\partial\hat{x}_{t+1}}{\partial\theta_{i}}$ $=\mathrm{Tr}\left[\mathcal{H}_{x}\frac{\partial U(\bm{\theta})}{\partial\theta_{i}}U_{enc}(\hat{X}_{t})\rho_{0}U_{enc}^{\dagger}(\hat{X}_{t})U^{\dagger}(\bm{\theta})\right]+\mathrm{Tr}\left[\mathcal{H}_{x}U(\bm{\theta})U_{enc}(\hat{X}_{t})\rho_{0}U_{enc}^{\dagger}(\hat{X}_{t})\frac{\partial U^{\dagger}(\bm{\theta})}{\partial\theta_{i}}\right]$ (7)
$\quad+\mathrm{Tr}\left[\mathcal{H}_{x}U(\bm{\theta})\frac{\partial U_{enc}(\hat{X}_{t})}{\partial\theta_{i}}\rho_{0}U_{enc}^{\dagger}(\hat{X}_{t})U^{\dagger}(\bm{\theta})\right]+\mathrm{Tr}\left[\mathcal{H}_{x}U(\bm{\theta})U_{enc}(\hat{X}_{t})\rho_{0}\frac{\partial U_{enc}^{\dagger}(\hat{X}_{t})}{\partial\theta_{i}}U^{\dagger}(\bm{\theta})\right].$
@@
Together with its conjugate term $\partial U_{enc}^{\dagger}(\hat{X}_{t})/\partial\theta_{i}$, each term in the last equation of Eq. (8) can be efficiently calculated using the parameter shift rule. This is because the coefficients can be evaluated with the current values of $m_{t}$, $\hat{x}_{t}$, and previous values of $\partial m_{t}/\partial\theta_{i}$, $\partial\hat{x}_{t}/\partial\theta_{i}$. Thus, while the gradient of $\hat{x}_{t+1}$ with respect to variable $\theta_{i}$ depends on all previous time steps, the gradient for a single time step can be computed in 6 evaluations based on the parameter shift rule, provided that the partial derivatives $\partial m_{t}/\partial\theta_{i}$, $\partial\hat{x}_{t}/\partial\theta_{i}$ are computed in order $t=1,2,\dots,L$. Since $m_{t+1}$ is obtained with the same quantum circuit as $\hat{x}_{t}$ but with a different measurement $\mathcal{H}_{m}$, computing $\partial m_{t}/\partial\theta_{i}$ will not increase the number of evaluations. The total complexity for computing $\partial L/\partial\theta_{i}$ is $\mathcal{O}(nL)$.
Our proposed model QDE avoids the barren plateau issue in quantum neural networks (QNNs) by utilizing multiple independent channels of QDE block with shallow circuit depths and weak entanglement. The shallow circuits limit the entanglement within each channel, preventing the volume-law scaling of entanglement that leads to barren plateaus [48]. Furthermore, the independence of channels reduces overall entanglement, decreasing the likelihood of encountering regions with vanishing gradients in the optimization landscape [49]. This design ensures a more navigable optimization landscape, enabling efficient and reliable training of the quantum neural network. Therefore, QDE’s structure effectively mitigates the barren plateaus in QNNs, providing a robust framework for quantum time series predictions [50].

### v4

@@ Abstract
The increasing focus on long-term time series prediction across various fields has been significantly strengthened by advancements in quantum computation. In this paper, we introduce a data-driven method designed for time series prediction with quantum dynamical embedding (QDE). This approach enables a trainable embedding of the data space into an extended state space, allowing for the recursive retrieval of time series information. Based on its independence of time series length, this method achieves depth-efficient quantum circuits that are crucial for near-term quantum computers. Numerical simulations demonstrate the model’s capability to predict not only wave signals but also more complex signals such as NARMA. Prediction accuracy improves with model scaling, and notably, the model achieves better accuracy on wave signal tasks with fewer parameters compared to QRC. Additionally, the model shows promising potential for denoising classical noise in wave signals, and when combined with error mitigation techniques for typical quantum noise, it enables reliable long-term prediction of wave signals. We implement this model, restricted to 2 qubits, on the Origin “Wukong” superconducting quantum processor as a simple proof-of-concept on NISQ devices. Furthermore, we provide theoretical analysis of the QDE’s dynamical properties for the 2-qubit case and discuss its potential universality. Overall, this study represents our first step towards leveraging near-term quantum devices for time series forecasting, offering insights into integrating data-driven learning with quantum dynamical embeddings.
@@ I Introduction
Figure 1: Circuits for QDE and Learning Protocols. (a) Illustration of a generic QDE architecture. The qubits for encoding the classical state are divided into two sets: the memory register (top, with $n_{m}$ qubits) for inputting memory $\bm{m}_{t}$, and the data register (bottom, with $n_{x}$ qubits) for inputting data $\bm{x}_{t}$ at time $t$. The circuit then undergoes evolution through a unitary operator $U(\bm{\theta})$ under a specified ansatz. Finally, selected observable ensembles are measured to extract classical information. With this architecture and the initial point $(\bm{m}_{0},\bm{x}_{0})$, the state pairs $(\bm{m}_{1},\bm{x}_{1}),(\bm{m}_{2},\bm{x}_{2}),(\bm{m}_{3},\bm{x}_{3}),\dots$ are obtained in an autoregressive way. (b) The protocol of QRC and QRNNs. The data is injected into the quantum system and measured to retrieve classical information. The memory register maintains the quantum state form and will lose coherence in the long run. (c) An enhanced QDE architecture with $M$ linear-composite channels. In this design, the QDE is available to explore the dynamics with higher degrees of freedom by either scaling the register qubit number or the composite channel number. The final signal $\bm{x}_{t}$ is retrieved through the superposition of sub-signals $\bm{x}_{t}^{1},\dots,\bm{x}_{t}^{M}$ from different channels. (d) Training and predicting protocols. In the training stage, initial data $\bm{x}_{0}$ is fed into the QDE block, and the output data $\hat{\bm{x}}_{i}$ is used to compute the cost function, with $i={1,\cdots,L}$ and time updated over $L$ steps. After the training, the system undergoes an additional $T$ steps of evolution to predict future values. Note that while the memory evolves simultaneously with the data in both stages, it does not contribute to the cost function calculation.
@@
Recent years have witnessed a surge in the application of time series prediction across a range of fields, including finance, physics, and engineering, underscoring its significance [1, 2, 3]. Among the various types of predictions, long-term prediction is especially pivotal as it entails forecasting future values of a series that extend beyond the duration of the available historical data. This means that the prediction horizon is longer than the time span of the training dataset, presenting a unique challenge in identifying and extrapolating underlying patterns and trends that may continue to evolve over time [4, 5]. The importance of long-term prediction is evident in strategic planning for climate change, economic forecasting, and technological forecasting, where decisions made today are based on projections that reach far into the future [6, 7]. Classical algorithms for time series prediction predominantly rely on long short-term memory (LSTM) networks [8, 9], echo state networks (ESNs) [10, 11], and reservoir computing (RC) [12]. These data-driven methods have shown potential in capturing dynamical features. Recent advances in quantum computing have demonstrated significant potential in the machine-learning domain [13, 14, 15, 16]. Quantum computers offer a promising avenue to overcome classical computational bottlenecks by enabling coherent processing of data within exponentially large Hilbert spaces, providing potential quantum advantages. Therefore, it is both intriguing and important to explore quantum-based approaches focused on time series processing, especially for realistic and challenging tasks such as long-term time series prediction. Recently, it has been demonstrated that classical data can be mapped into a quantum feature space using certain embeddings [14, 15, 17, 18], which is further utilized to approximate a variety of functions. These universal approximation properties [13, 19, 20, 21] facilitate the construction of inference functions for modeling temporal data patterns, suggesting the potential of quantum systems to effectively model the complex dynamics inherent in long-term time series.
@@
Current quantum algorithms for time series processing, such as quantum reservoir computing (QRC) [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], quantum recurrent neural networks (QRNNs) [35, 36, 37], and quantum dynamic mode decomposition (QDMD) [38], have demonstrated significant potential in understanding temporal dependencies in sequential data. However, in QRC and QRNNs, the circuit depth is determined by a constant factor ($C$) from the ansatz, multiplied by the time sequence length ($L$), i.e., $C\times L$, which poses challenges for noisy intermediate-scale quantum (NISQ) devices. The increasing circuit depth threatens the maintenance of quantum coherence over time, which is crucial for the accuracy of long-term predictions. Although the QRC model has shown significant performance of time series prediction, its requirement for deep circuits to maintain temporal dependencies makes it impractical for NISQ devices due to exacerbated decoherence and noise, which undermine the performance of quantum algorithms in realistic settings. Efficiently utilizing these devices for complex tasks like long-term time series prediction without deep circuits is essential for advancing practical quantum computing applications. Therefore, there is a pressing need for methods that can predict long-term time series with shallow quantum circuits, addressing the limitations posed by current quantum hardware. Solutions have been explored through both theoretical and physical implementations on various platforms [39, 40, 41, 42, 43, 44, 45], yet it remains an open question.
@@
In this study, we present a data-driven approach for time series prediction using quantum dynamical embedding (QDE). Our approach allows for the dynamic embedding of original data into an extended state space and facilitates the data-driven learning of embedding via trainable quantum-gate parameters, i.e., in a quantum-neural-network style as in [36]. It differs from feedback-based methods [46, 47, 34] in the placement and optimization of tunable parameters: the parameters reside within the circuit ansatz and optimized iteratively, rather than learned by regression at the readout as in [22]. Specifically, we recurrently retrieve time series data from measurements of the QDE at the current time step. This design splits the long-term time sequence into independent input-prediction pairs, ensuring that for each prediction task the qubit count and circuit depth $C$ are determined by the ansatz architecture, irrespective of the time series length. Although the circuit depth is reduced, a full prediction of the time sequence is achieved by repeating the circuit $L$ times with different inputs. Consequently, it mitigates issues related to decoherence and error accumulation over time.
@@
We demonstrate the effectiveness of our method through numerical experiments on wave signal prediction as well as more complex tasks like NARMA. Additionally, we evaluate its denoising capabilities on the same prediction tasks when exposed to random noise, further showcasing the QDE’s versatility. Notably, a comparative analysis with existing methodologies reveals that our approach offers significant advantages in terms of average prediction accuracy and computational efficiency on the same tasks. We also recognize limitations related to the training procedure, such as the presence of barren plateau, local minima and instability, which point to important directions for further model optimization.
To validate the noise resistance of our method, we perform long-term predictions using a noisy simulator with a learnable error-cancellation layer (LECL) following the QDE. Additionally, we extend our implementation to the Origin “Wukong” superconducting quantum computer, integrating an LECL to generate time series for proof-of-concept demonstration. This demonstrates the potential of leveraging current quantum computing capabilities for real-world time series prediction challenges.
Theoretically, we explore the dynamics of a specific 2-qubit case, providing insights into the connection between the QDE and the inherent characteristics of time series. Developing a more general theoretical framework, together with demonstrating performance improvements for broader time series processing tasks, remains an open challenge. Although this approach does not fully replace QRC, the data embedding method offers a promising and insightful direction for quantum machine learning applied to time series analysis.
@@ II.1 The quantum dynamical embedding model
is the encoded input state. $U(\bm{\theta})$ is a trainable parameterized quantum circuit. The encoding strategies $U_{\text{enc}}$ may vary; a commonly used scheme is introduced later in Sec. II.4. Our method relies on the assumption that the embedding effectively maps the original non-Markovian dynamics to an (approximate) Markovian form. Accordingly, in the present work, we neglect time-delayed memory terms, denoted $\bm{m}_{t-i}$ for $i>0$. The memory and data of the next step are extracted from observing of their corresponding subspaces, namely
@@
where $\mathcal{H}_{m}$ and $\mathcal{H}_{x}$ are the observable vector with length $n_{m}$ and $n_{x}$, respectively. In the following, we explicitly consider the Pauli-$\sigma_{z}$ operator as the observable. This means that $(\mathcal{H}_{m},\mathcal{H}_{x})$ correspond to measuring $\sigma_{z}$ on each qubit individually.
@@
The QDE distinguishes itself from previous studies of QRC and QRNNs in the manner of memory information storage. Unlike these models, where the quantum system serves as a memory reservoir for historical data, our approach overcomes the issue of coherence loss—a challenge in long-term data embedding and retrieval—through the QDE model, as illustrated in Fig. 1(b). Recent works [46, 34] also adopt a reinjection-based approach to memory, but differ from QDE in their memory-state reconstruction procedures and in the overall learning and optimization framework.
@@ II.2 Training and predicting
Building on the findings of previous research [22, 36], the embedding of time series and the evolving of quantum systems require a circuit depth of $\mathcal{O}(L)$. Comparatively, our approach marks a notable advancement by achieving this with a significantly reduced circuit depth of $\mathcal{O}(\Gamma)$, where $\Gamma$ represents the base depth of a single QDE circuit block.
@@ II.3 Gradient evaluation
Referring to the definition of $\hat{x}_{t+1}$ in Eq. (3) and the explicit expression of $\mathcal{H}_{m}=\sigma_{z}^{0}$ and $\mathcal{H}_{x}=\sigma_{z}^{1}$, this expression can be equivalently written as:
$\frac{\partial\hat{x}_{t+1}}{\partial\theta_{i}}$ $=\frac{\partial}{\partial\theta_{i}}\operatorname{Tr}\left(\sigma_{z}^{1}|\psi_{t}\rangle\langle\psi_{t}|\right).$ (6)
$=\frac{\partial}{\partial\theta_{i}}\mathrm{Tr}\left[\sigma_{z}^{1}U(\bm{\theta})U_{enc}(\hat{X}_{t})\rho_{0}U_{enc}^{\dagger}(\hat{X}_{t})U^{\dagger}(\bm{\theta})\right],$
@@
$\frac{\partial\hat{x}_{t+1}}{\partial\theta_{i}}$ $=\mathrm{Tr}\left[\sigma_{z}^{1}\frac{\partial U(\bm{\theta})}{\partial\theta_{i}}U_{enc}(\hat{X}_{t})\rho_{0}U_{enc}^{\dagger}(\hat{X}_{t})U^{\dagger}(\bm{\theta})\right]+\mathrm{Tr}\left[\sigma_{z}^{1}U(\bm{\theta})U_{enc}(\hat{X}_{t})\rho_{0}U_{enc}^{\dagger}(\hat{X}_{t})\frac{\partial U^{\dagger}(\bm{\theta})}{\partial\theta_{i}}\right]$ (7)
$\quad+\mathrm{Tr}\left[\sigma_{z}^{1}U(\bm{\theta})\frac{\partial U_{enc}(\hat{X}_{t})}{\partial\theta_{i}}\rho_{0}U_{enc}^{\dagger}(\hat{X}_{t})U^{\dagger}(\bm{\theta})\right]+\mathrm{Tr}\left[\sigma_{z}^{1}U(\bm{\theta})U_{enc}(\hat{X}_{t})\rho_{0}\frac{\partial U_{enc}^{\dagger}(\hat{X}_{t})}{\partial\theta_{i}}U^{\dagger}(\bm{\theta})\right].$
@@
Together with its conjugate term $\partial U_{enc}^{\dagger}(\hat{X}_{t})/\partial\theta_{i}$, each term in the last equation of Eq. (8) can be efficiently calculated using the parameter shift rule. This is because the coefficients can be evaluated with the current values of $m_{t}$, $\hat{x}_{t}$, and previous values of $\partial m_{t}/\partial\theta_{i}$, $\partial\hat{x}_{t}/\partial\theta_{i}$. Thus, while the gradient of $\hat{x}_{t+1}$ with respect to variable $\theta_{i}$ depends on all previous time steps, the gradient for a single time step can be computed in 6 evaluations based on the parameter shift rule, provided that the partial derivatives $\partial m_{t}/\partial\theta_{i}$, $\partial\hat{x}_{t}/\partial\theta_{i}$ are computed in order $t=1,2,\dots,L$. Since $m_{t+1}$ is obtained with the same quantum circuit as $\hat{x}_{t}$ but with a different measurement $\mathcal{H}_{m}$, computing $\partial m_{t}/\partial\theta_{i}$ will not increase the number of evaluations. Taking into account the sampling error for each expectation value $\langle\sigma_{z}\rangle$, the total sample complexity for estimating the gradient $\partial\mathcal{L}_{mse}/\partial\bm{\theta}$ scales as $\mathcal{O}(\frac{4(n+1)L|\{\theta_{i}\}|}{\varepsilon^{2}})$, where $n$ is the number of qubits, $L$ is the sequence length, $|\{\theta_{i}\}|$ is the number of circuit parameters, and $\varepsilon$ denotes the desired precision for observable measurements.
As a specialized type of quantum neural network (QNN), it is natural to consider whether the QDE model suffers from the barren plateau phenomenon, where the loss function or its gradients concentrate and vanish as the system size increases. This issue has been extensively studied, with analyses of its origins and methods to mitigate or avoid it [52, 53, 54, 55, 56, 57]. Since the QDE is constructed from the QNN building blocks, it is important to evaluate the impact of barren plateaus on its trainability. In Section III.1, we provide numerical results assessing the gradient concentration of the QDE model to better understand its behavior in this regard.